A 107 g sample of CH3CH2Br (Bromoethane) gas has a volume of 4.375 L at a temperature of 36°C.
i. Calculate the pressure using the Ideal Gas Law.
ii. Calculate the corrected pressure using the Van der Waals Equation.
atm×L2 L a=11.723 mol2 b=0.08406 mol
1
Expert's answer
2020-04-15T13:59:04-0400


Molar mass of Bromoethane is 109 gms.

Thus, number of moles = "n=107\/109"

Given : "V=4.375 L ; T=36\u00b0C=273+36=309K"

We know Gas constant = "R=0.0821 Latm mol^{-1}K^{-1}"


i. Ideal gas law : "PV=nRT \implies P=nRT/V"

"P=107*0.0821*309\/(109*4.375)=5.692 atm" ---(Answer)


ii. Van der Waals Equation : "(P+an^2\/V^2)(V-nb)=nRT"

Given : "a=11.723atmL^2 mol^{-2}; b=0.08406L mol^{-1}"

Thus, "(P+(11.723*107^2\/(109^2*4.375)))(4.375-""(107*0.08406\/109)=107*0.0821*309\/109"

"\implies (P+2.582)(4.375-0.0825)=24.9"

"\implies P=(24.9/4.292)-2.582=3.22 atm" ----(Answer)